OpenStudy (anonymous):

Solving a definite integral with u-substitution: [(x^4 + x)^5](4x^3 + 1) dx from 0 to 1.

OpenStudy (anonymous):

\[\int_0^1(x^4+x)(4x^3+1)dx\] put \(u=x^4+x, du=(4x^3+1)dx,u(0)=0, u(1)=2\) and change to \[\int_0^2udu\]

OpenStudy (anonymous):

oh i missed the exponent of \(5\) sorry, it would be \[\int_0^2u^5du\] in that case

OpenStudy (dan815):

noooo

OpenStudy (dan815):

i wanted to do it

OpenStudy (anonymous):

Thanks for your quick reply, I'll give it a go.

OpenStudy (anonymous):

yw

OpenStudy (dan815):